Find resistance of a conductor with increasing crosssection area.  The radius of one end is a and another end is b.  Find in terms of rho,  l, a, b

Grade:12

1 Answers

Vikas TU
6 years ago
As we as a whole realize that 
Resistance = rho*l/Crosssectional zone 
Zone = pi*(a^2-b^2) 
Resistance = (rho*l)/( pi*(a^2-b^2))